### v1

# Cardinalities in proofreading

Abstract
In 1975, G. Pólya pointed out that if two proofreaders found $a$, resp. $b$, errors in a text, and both found $c$ errors, then the unknown number $e$ of all errors is approximated by $e\approx\frac{ab}{c}$. We devise a realistic combinatorial model of Pólya’s result and estimate its performance by the Probabilistic Method.
1 Introduction. The $(p,n)$-proofreader
In [8], see also [1, 3, 4, 7], G. Pólya devised a conjuring trick. Suppose that a proofreader finds $a$ errors in a text, that another proofreader finds $b$ errors, and that the number of errors detected by both is $c$. How can one guess, with these data, the unknown number $e$ ($>0$) of all errors? Well, $a\approx pe$ and $b\approx qe$, where $p$ and $q$ ($\in(0,1]$) are probabilities of error detection for the proofreaders. Also, $c\approx pqe$ due to their independence. Then
$\frac{ab}{c}\approx\frac{peqe}{pqe}=e\,.$
Pólya’s trick, brilliant as it is, is not a rigorous mathematical theorem. In our article, we elaborate it in a combinatorial theorem. We state this Theorem 1.3 below and prove it in the next section. The last third section contains comments on probabilistic vs. discrete models. For $n$ in $\mathbb{N}$ (=$=\{1,2,\dots\}$) we set $[n]=\{1,2,\dots,n\}$. Let $B_{n}=\{0,1\}^{n}$ be the set of binary $n$-tuples
[…]
We denote by $\|\overline{b}\|=\sum_{i=1}^{n}b_{i}$ the number of $1$s in $\overline{b}$. For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1,\dots\}$) its cardinality (number of elements), for example $|B_{n}|=2^{n}$. If $\overline{s}$ is a finite tuple, then $|\overline{s}|$ denotes the number of terms in it. By $\mathbb{Q}$ we denote the set of fractions.
Let $n\in\mathbb{N}$ and $p=\frac{k}{l}\in(0,1]\cap\mathbb{Q}$, with $k,l\in\mathbb{N}$, $k\leq l$ and coprime $k,l$. We always assume that $p>0$. The $(p,n)$-proofreader is the pair $(P,\overline{s})$ in which
$P=P(\overline{a},\,\overline{b})\colon B_{n}\times B_{n}\to B_{n},\ P(\overline{a},\,\overline{b})=(a_{1}b_{1},\,a_{2}b_{2},\,\dots,\,a_{n}b_{n})\,,$
and
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N}),\ N=l^{n}\,\text{ and }\,\overline{s}_{j}\in B_{n}\,,$
is a uniquely determined $N$-tuple defined as follows. We take the map $f\colon[l]\to\{0,1\}$, $f(x)=1$ for $x\in[k]$ and $f(x)=0$ for $x\in[l]\setminus[k]$, and extend it to
$f\colon[l]^{n}\to B_{n},\ f((a_{1},\,a_{2},\,\dots,\,a_{n}))=(f(a_{1}),\,f(a_{2}),\,\dots,\,f(a_{n}))\,.$
Then we take the lexicographic ordering
$(\overline{u}_{1},\,\overline{u}_{2},\,\dots,\,\overline{u}_{N})$
of $[l]^{n}$, so that $\overline{u}_{1}=(1,1,\dots,1)$, $\overline{u}_{2}=(1,1,\dots,1,2)$, $\dots$, $\overline{u}_{N}=(l,l,\dots,l)$. Any ordering of $[l]^{n}$ would do, but we want that $(p,n)$-proofreaders be unique objects. Finally, we set
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})=(f(\overline{u}_{1}),\,f(\overline{u}_{2}),\,\dots,\,f(\overline{u}_{N}))\,.$
Let $(P,\overline{s})$ be the $(p,n)$-proofreader and $\overline{t}\in B_{n}$ be a text, in which the errors are positions of $1$s. Then for $j\in[N]=[l^{n}]$, where $l$ is the denominator of $p$, the proofreader finds in the text $\overline{t}$ in the $j$-th run $\overline{s}_{j}$ exactly $\|P(\overline{s}_{j},\overline{t})\|$ errors. We show that proofreaders work in the desired way.
Proposition 1.1
Let $(P,\overline{s})$, where
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\ \ (\in B_{n}^{N})\,,$
be the $(p,n)$-proofreader. Then for every set $I\subset[n]$ we have
$N^{-1}\cdot\big{|}\big{\{}j\in[N]\colon\;i\in I\Rightarrow s_{j,\,i}=1\big{\}}\big{|}=p^{|I|}\,.$
Proof. Let $\overline{s}$ and $I$ be as stated. It is easy to see from the definition of $\overline{s}$ that the number of $j\in[N]$ such that $s_{j,i}=1$ for every $i\in I$ is $k^{|I|}\cdot l^{n-|I|}$. But then
$N^{-1}\cdot k^{|I|}\cdot l^{n-|I|}=\frac{k^{|I|}l^{n-|I|}}{l^{n}}=\bigg{(}\frac{k}{l}\bigg{)}^{|I|}=p^{|I|}\,.$
$\Box$
Proofreading of a text is usually done only once, or a few times. We think of the $N$ runs $\overline{s}_{j}$ as of performances of the proofreader in possible worlds ([9]).
We need two independent proofreaders. Let $n\in\mathbb{N}$, $p,q\in(0,1]\cap\mathbb{Q}$ and let $(P,\overline{s})$ and $(P,\overline{u})$ be the $(p,n)$-proofreader and the $(q,n)$-proofreader, respectively, with $|\overline{s}|=M$ and $|\overline{u}|=N$. Two (independent) $(p,q,n)$-proofreaders is the pair $(P,\overline{S})$ such that
$\overline{S}$ $=$ $(\overline{S}_{1},\,\overline{S}_{2},\,\dots,\,\overline{S}_{MN})$
$=$ $((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{1},\,\overline{u}_{N}),\,(\overline{s}_{2},\,\overline{u}_{1}),\,(\overline{s}_{2},\,\overline{u}_{2}),\,\dots$
$\dots,\,(\overline{s}_{2},\,\overline{u}_{N}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{1}),\,(\overline{s}_{M},\,\overline{u}_{2})\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))$
($\in(B_{n}^{2})^{MN}$). Again, any ordering of the $MN$ pairs $(\overline{s}_{j},\overline{u}_{j^{\prime}})$ would do, but we want that pairs of proofreaders be unique. For any $j\in[MN]$ we write
$\overline{S}_{j}=\big{(}\overline{s}_{v_{j}},\,\overline{u}_{w_{j}}\big{)}\,.$
We show that two proofreaders work in the desired way.
Proposition 1.2
Let $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
be two $(p,q,n)$-proofreaders. The following holds.
  1. 1.
    For every set $I\subset[n]$ we have
    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow s_{v_{j},\,i}=1\big{\}}\big{|}=p^{|I|}\,.}$
  2. 2.
    For every set $I\subset[n]$ we have
    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow u_{w_{j},\,i}=1\big{\}}\big{|}=q^{|I|}\,.}$
  3. 3.
    For every set $I\subset[n]$ we have
    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow s_{v_{j},\,i}=u_{w_{j},\,i}=1\big{\}}\big{|}=(pq)^{|I|}\,.}$
Proof. Let $\overline{S}$ and $I$ be as stated. 1. Let $r$ be the number of $j\in[M]$ such that $s_{j,i}=1$ for every $i\in I$. By Proposition 1.1, $\frac{r}{M}=p^{|I|}$. Thus the stated product equals
${\textstyle\frac{1}{MN}\cdot rN=\frac{r}{M}=p^{|I|}\,.}$
2. The argument is symmetric, with the number $s$ of $j^{\prime}\in[N]$ such that $u_{j^{\prime},i}=1$ for every $i\in I$. 3. Now the stated product equals
${\textstyle\frac{1}{MN}\cdot rs=\frac{r}{M}\cdot\frac{s}{N}=p^{|I|}\cdot q^{|I|}=(pq)^{|I|}\,.}$
$\Box$
We state our main result, a combinatorial version of Pólya’s trick.
Theorem 1.3
Let $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
be two $(p,q,n)$-proofreaders and let
${\textstyle c(p,\,q)=\frac{1}{2}\big{(}\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big{)}^{-1}\,.}$
Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have for every number $a\in[0,m]$ that
${\textstyle\frac{1}{MN}}\cdot\big{|}\big{\{}j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq a\big{\}}\big{|}\geq$
${\textstyle\geq 1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})\,.}$
For $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$ we define the fraction as +$\infty$+ ∞.
Pólya’s estimate has distance at most $a$ from the number of errors for at least the fraction $1-\dots$ of pairs $(\overline{s}_{v_{j}},\overline{u}_{w_{j}})$ of runs of the two proofreaders. Next, in one extreme regime of Theorem 1.3 we minimize the bound $a$ on the distance. In the other extreme we maximize the fraction of pairs of runs for which the distance is still reasonably small. Deduction of both corollaries is immediate.
Corollary 1.4
In the setup of Theorem 1.3, if $\omega(m)\colon\mathbb{N}\to\mathbb{N}$ is any function such that $\omega(m)\leq\sqrt{m}$ and $\omega(m)\to+\infty$, then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}}\cdot\big{|}\big{\{}j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq\omega(m)\sqrt{m}\big{\}}\big{|}\geq$
$\geq 1-6\exp(-2c(p,\,q)^{2}\omega(m)^{2})\to 1\ \ (m\to+\infty)\,.$
Corollary 1.5
In the setup of Theorem 1.3, if $\varepsilon\in(0,1)$ then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}}\cdot\big{|}\big{\{}j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq\varepsilon m\big{\}}\big{|}\geq$
$\geq 1-6\exp(-2\varepsilon^{2}c(p,\,q)^{2}m)\to 1\ \ (m\to+\infty)\,.$
We prove Theorem 1.3 next.
2 Proof of Theorem 1.3
We deduce Theorem 1.3 from two propositions.
Proposition 2.1
Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\big{|}\|P(\overline{s}_{v_{j}},\,\overline{t})\|-pm\big{|}\leq a\big{\}}\big{|}\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
An analogous bound holds when $\overline{s}_{v_{j}}$ replaced by $\overline{u}_{w_{j}}$, and $pm$ by $qm$.
Proposition 2.2
Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\big{|}\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm\big{|}\leq a\big{\}}\big{|}\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
We prove both propositions at the conclusion of this section. Now we deduce from them Theorem 1.3.
Deduction of Theorem 1.3 from Propositions 2.1 and 2.2. Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
are two $(p,q,n)$-proofreaders, that $\overline{t}\in B_{n}$ is a text with $\|\overline{t}\|=m\geq 1$ errors and that $a\in[0,m]$. We set
${\textstyle b=c(p,\,q)\cdot a=\frac{1}{2}\big{(}\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big{)}^{-1}\cdot a\ \ (\leq\frac{pqm}{2})\,.}$
For $j\in[MN]$, let $A$ be the set of $j$ such that $|\,\|P(\overline{s}_{v_{j}},\overline{t})\|-pm|\leq b$, $B$ be the set of $j$ such that $|\,\|P(\overline{u}_{w_{j}},\overline{t})\|-qm|\leq b$, $C$ be the set of $j$ such that
$|\,\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm|\leq b$
and, finally, let $D$ be the set we are interested in, the set of $j$ such that
$|R-m|:=\big{|}\,\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|\cdot\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|^{-1}-m\big{|}\leq a$
($\|\dots\|\cdot\|\dots\|\cdot\|\dots\|^{-1}=+\infty$ if $\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=0$). Thus if $j\in A$, resp. $j\in B$, then
[…]
where $|\delta_{1}|\leq\frac{b}{pm}$, resp. $|\delta_{2}|\leq\frac{b}{qm}$. Similarly, if $j\in C$ then
${\textstyle\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=(1+\delta_{3})pqm,\ |\delta_{3}|\leq\frac{b}{pqm}\leq\frac{1}{2}\,.}$
If $j\in A\cap B\cap C$ then
$|R-m|$ $=$ ${\textstyle\big{|}\frac{(1+\delta_{1})(1+\delta_{2})}{1+\delta_{3}}m-m\big{|}}$
$=$ ${\textstyle m\left|(1+\delta_{1}+\delta_{2}+\delta_{1}\delta_{2})\sum_{k=0}^{\infty}(-1)^{k}\delta_{3}^{k}-1\right|}$
$=$ ${\textstyle m\left|\sum_{k=1}^{\infty}(-1)^{k}\delta_{3}^{k}+(\delta_{1}+\delta_{2}+\delta_{1}\delta_{2})\sum_{k=0}^{\infty}(-1)^{k}\delta_{3}^{k}\right|}$
$\leq$ $m\left(2|\delta_{3}|+2(|\delta_{1}|+|\delta_{2}|+|\delta_{1}|\cdot|\delta_{2}|)\right)$
$\leq$ $2b{\textstyle\big{(}\frac{1}{pq}+\frac{1}{p}+\frac{1}{q}+\frac{1}{pq}\big{)}=a\,,}$
and hence $D\supset A\cap B\cap C$. By Propositions 2.1 and 2.2,
${\textstyle\frac{1}{MN}\cdot|D|\geq 1-\frac{|A^{c}|+|B^{c}|+|C^{c}|}{MN}\geq 1-6\exp(-\frac{2b^{2}}{m})=1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})\,.}$
$\Box$
We develop tools for proofs of Propositions 2.1 and 2.2. Let $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
be two $(p,q,n)$-proofreaders. We consider the finite uniform probability space
$P=P(MN)=([MN],\,\mathcal{P}([MN]),\,\mathrm{Pr})\,,$
in which $\mathrm{Pr}(\{j\})=\frac{1}{MN}$, and for $i\in[n]$ the three $n$-tuples of $P$-events
$A_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=1\},\ A_{i}^{\prime}=\{j\in[MN]\colon\;u_{w_{j},i}=1\}$
and
$C_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=u_{w_{j},i}=1\}\,.$
By Propositions 1.1 and 1.2, each of these three $n$-tuples is independent. Further, $\mathrm{Pr}(A_{i})=p$, $\mathrm{Pr}(A_{i}^{\prime})=q$ and $\mathrm{Pr}(C_{i})=pq$ for every $i\in[n]$. Recall that $B^{c}$ denotes the complementary event to the event $B$ in a probability space. For example,
$A_{i}^{c}=[MN]\setminus A_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=0\}\,.$
Let $X\subset[n]$ be a nonempty set. For any set $Y\subset X$ we consider the three $P$-events
${\textstyle A_{Y,X}=\bigcap_{i\in Y}A_{i}\cap\bigcap_{i\in X\setminus Y}A_{i}^{c},\ A_{Y,X}^{\prime}=\bigcap_{i\in Y}A_{i}^{\prime}\cap\bigcap_{i\in X\setminus Y}(A_{i}^{\prime})^{c}}$
and
${\textstyle C_{Y,X}=\bigcap_{i\in Y}C_{i}\cap\bigcap_{i\in X\setminus Y}C_{i}^{c}\,.}$
We use the following key theorem which restates [2, Corollary A.1.7 on p. 325]. It is proven on [2, pp. 321–325].
Theorem 2.3
Let $n\in\mathbb{N}$, $X\subset[n]$ with $|X|=m\geq 1$ and let $a>0$ be a real number. With the above notation we have the bound
${\textstyle\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
Two analogous bounds hold when $A_{Y,X}$ is replaced by $A_{Y,X}^{\prime}$, resp. $C_{Y,X}$, and $pm$ by $qm$, resp. $pqm$.
We explain in detail this restatement. Let $m\in\mathbb{N}$, $p\in[0,1]$, $X_{i}$ for $i\in[m]$ be $m$ independent two-valued real random variables in a probability space, with the common distribution $\mathrm{Pr}(X_{i}=1-p)=p$ and $\mathrm{Pr}(X_{i}=-p)=1-p$, and let
$X_{0}=X_{1}+X_{2}+\dots+X_{m}\,.$
The random variable $X_{0}$ has the centered binomial distribution $B(m,p)-pm$. The estimate [2, Corollary A.1.7], more precisely a particular case of it with all $p_{i}=p$, says the following.
Theorem 2.4 (Corollary A.1.7 in [2])
For every real number $a>0$,
${\textstyle\mathrm{Pr}(|X_{0}|>a)<\exp(-\frac{2a^{2}}{m})\,.}$
What does the independence of random variables $X_{i}$, $i\in[m]$, which is assumed in Theorem 2.4, precisely mean? Some events $C_{1}$, $C_{2}$, $\dots$, $C_{m}$ in a probability space are independent, if for every set $I\subset[m]$ we have
${\textstyle\mathrm{Pr}\big{(}\bigcap_{i\in I}C_{i}\big{)}=\prod_{i\in I}\mathrm{Pr}(C_{i})\,.}$
These events are strongly independent, if every $m$-tuple of events $D_{1}$, $D_{2}$, $\dots$, $D_{m}$, where $D_{i}\in\{C_{i},C_{i}^{c}\}$, is independent. A well known lemma, whose proof we leave for the interested reader as an exercise, says that any independent events are strongly independent. Discrete (i.e., with at most countable images) real random variables $Y_{1}$, $Y_{2}$, $\dots$, $Y_{m}$ are independent if for every $m$-tuple
$\overline{a}=(a_{1},\,a_{2},\,\dots,\,a_{m})\in\mathbb{R}^{m}$
the events $\mathrm{Pr}(Y_{i}=a_{i})$, $i\in[m]$, are independent. If the $Y_{i}$ are two-valued, the image of $Y_{i}$ being $\{a_{i},b_{i}\}$ with $a_{i}\neq b_{i}$, then, due to the mentioned lemma, the random variables $Y_{i}$ are independent if for any single $m$-tuple $\overline{c}\in\mathbb{R}^{m}$ with $c_{i}\in\{a_{i},b_{i}\}$ the events $\mathrm{Pr}(Y_{i}=c_{i})$, $i\in[m]$, are independent.
We deduce Theorem 2.3 from Theorem 2.4. We consider only events $A_{Y,X}$, for events $A_{Y,X}^{\prime}$ and $C_{Y,X}$ we would argue similarly, relying on part 2 and 3 of Proposition 2.2. So let $(P,\overline{S})$, with
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
be as above two $(p,q,n)$-proofreaders and let $X\subset[n]$ with $|X|=m\geq 1$. We may assume that $X=[m]$. For $i\in[m]$ we define two-valued $P(MN)$-random variables $X_{i}$ by setting $X_{i}(j)=1-p$ if $j\in A_{i}$, and $X_{i}(j)=-p$ if $j\in[MN]\setminus A_{i}$. We set $X_{0}=\sum_{i=1}^{m}X_{i}$. Then for any subset $Y\subset X$ we have
$\mathrm{Pr}(A_{Y,X})=\mathrm{Pr}(X_{0}=|Y|-pm)\,.$
By part 1 of Proposition 2.2 and the above discussion of independence, the random variables $X_{i}$, $i\in[m]$, are independent. Let $a>0$ be a real number. By applying Theorem 2.4 to the random variable $X_{0}$ we get the estimate in Theorem 2.3.
We proceed to proofs of Propositions 2.1 and 2.2.
Proof of Proposition 2.1. Suppose that $(P,\overline{S})$, where
$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Let $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$. We may assume that the $1$s of $\overline{t}$ are on the first $m$ positions. We consider the $P(MN)$-events $A_{i}$, $i\in[m]$. Let $Y\subset[m]=X$. Then
${\textstyle\frac{1}{MN}|\{j\in[MN]\colon\;\text{for $i\in[m]$ we have $s_{v_{j},i}=1$ iff $i\in Y$}\}|=\mathrm{Pr}(A_{Y,X})\,.}$
Thus for any $k\in\mathbb{N}_{0}$ we have
${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\|P(\overline{s}_{v_{j}},\,\overline{t})\|=k\big{\}}\big{|}=\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |Y|=k\end{subarray}}\mathrm{Pr}(A_{Y,X})\,.}$
The stated estimate therefore follows from the first part of Theorem 2.3. The second part of Proposition 2.1 follows by a very similar argument from the second part of Theorem 2.3. $\Box$
Proof of Proposition 2.2. Let $(P,\overline{S})$, $a>0$ and $\overline{t}$ be as in the previous proof. We consider the $P(MN)$-events $C_{i}$, $i\in[m]$, and argue as in the previous proof, relying on the third part of Theorem 2.3. $\Box$
3 Concluding remarks
By [5, p. 109], “The subject of probability is wholly free from real paradoxes.” This is not precise — not only books like [10] are devoted to paradoxes in probability theory, but some of these paradoxes are truly real. At their roots is the unrealistic nature, or even total absence, of physical experiments that should be modeled by the probabilistic model, or PM, at hand. In these paradoxical cases we often have a PM that does not model anything realistic. For example, take well known Bertrand’s paradox which asks about the probability that random chord in the unit circle in the plane has length at least $\sqrt{3}$ (the side of the inscribed equilateral triangle). The noun phrase “random chord” is problematic because it does not refer to anything clearly defined. In contrast, noun phrases like “random (finite!) graph” or “random subset (of a finite set)” are unproblematic; these concepts are easily modeled by finite uniform probability spaces and events in them, like the above $P(MN)$ and the $P(MN)$-events $A_{i}$, $A_{i}^{\prime}$ and $C_{i}$, $i\in[n]$. Mathematical chords have zero width and form an uncountable set. People try to resolve Bertrand’s paradox by rushing in defining PM on the set of mathematical chords. Instead, they should step back and consider physical chords with width $w>0$. A realistic and convincing physical discrete model, or DM, for selection of random physical chords should be devised. Only when such DM is understood, we can send $w$ to 00 and see what happens. This approach is implemented in [6]. We hope to rework [6] in near future in a better form.
In this article we approached Pólya’s proofreaders from the DM perspective. Returning to Proposition 1.1, we see that if $(P,\overline{s})$, where
$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\,,$
is the $(p,n)$-proofreader, then in the uniform probability space
$P(N)=([N],\,\mathcal{P}([N]),\,\mathrm{Pr})\,,$
with $\mathrm{Pr}(\{j\})=\frac{1}{N}$, the $n$ $P(N)$-events
$A_{i}=\{j\in[N]\colon\;s_{j,i}=1\},\ i\in[n]\,,$
are independent and model random subset of $[n]$ with probability of an element equal to $p$. However, our construction requires $p$ ($\in(0,1]$) be a rational number. This is just as right because we cannot have a realistic physical experiment with probabilities like $\frac{1}{\sqrt{2}}$. Thus we conclude our article with the comment, which may come as a surprise to some, that the random graph $G(n,p)$ with probability of an edge equal to an irrational number $p$ is not really a realistic model.

### v2

@@
 In 1975, G.
-Pólya pointed out that if two proofreaders found $a$, resp.
-$b$, errors in a text, and both found $c$ errors, then the unknown number $e$ of all errors is approximated by $e\approx\frac{ab}{c}$.
-We devise a realistic combinatorial model of Pólya’s result and estimate its performance by the Probabilistic Method.
-1 Introduction. The $(p,n)$-proofreader
-In [8], see also [1, 3, 4, 7], G.
-Pólya devised a conjuring trick.
-Suppose that a proofreader finds $a$ errors in a text, that another proofreader finds $b$ errors, and that the number of errors detected by both is $c$.
-How can one guess, with these data, the unknown number $e$ ($>0$) of all errors?
-Well, $a\approx pe$ and $b\approx qe$, where $p$ and $q$ ($\in(0,1]$) are probabilities of error detection for the proofreaders.
-Also, $c\approx pqe$ due to their independence.
-Then
-$\frac{ab}{c}\approx\frac{peqe}{pqe}=e\,.$
-Pólya’s trick, brilliant as it is, is not a rigorous mathematical theorem.
-In our article, we elaborate it in a combinatorial theorem.
-We state this Theorem 1.3 below and prove it in the next section.
-The last third section contains comments on probabilistic vs.
-discrete models.
-For $n$ in $\mathbb{N}$ (=$=\{1,2,\dots\}$) we set $[n]=\{1,2,\dots,n\}$.
-Let $B_{n}=\{0,1\}^{n}$ be the set of binary $n$-tuples
-[…]
-We denote by $\|\overline{b}\|=\sum_{i=1}^{n}b_{i}$ the number of $1$s in $\overline{b}$.
-For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1,\dots\}$) its cardinality (number of elements), for example $|B_{n}|=2^{n}$.
-If $\overline{s}$ is a finite tuple, then $|\overline{s}|$ denotes the number of terms in it.
+Pólya suggested that if two proofreaders found $a$ and $b$ errors in a text, of which $c$ errors were found by both of them, then a reasonable approximation of the unknown number $e$ of all errors is $e\approx ab/c$.
+We justify this formula by constructing a realistic model of proofreaders and estimating the efficiency of this model with the help of the bound on large deviations in the Probabilistic Method.
+In conclusion we discuss the distinction between realistic and probabilistic models of problems.
+Dedicated to the memory of my mother Blanka Klazarová (1937–2026)
+1 Introduction. $(p,n)$-proofreaders
+In [7], G.
+Pólya considered the following problem (see also [1, 3, 4, 6]).
+Suppose that a proofreader found $a$ errors in a text, that another proofreader found $b$ errors, and that the number of errors detected by both of them is $c$.
+With these data, how could one approximate the unknown number $e$ of all errors in the text?
+Pólya’s solution is this.
+We have $a\approx pe$ and $b\approx qe$, where $p$ and $q$ in $(0,1]$ are the probabilities of error detection for the two proofreaders.
+Their independence gives $c\approx pqe$.
+Thus we get the approximation
+$e=\frac{peqe}{pqe}\approx\frac{ab}{c}\,.$
+Pólya’s interesting argument is not supported by any rigorous theorem.
+In this article we provide one.
+It is Theorem 1.5 which we prove in Section 2.
+In Section 3 we discuss the two kinds of mathematical models one can devise to solve problems: realistic models and probabilistic models.
+For $n$ in $\mathbb{N}=\{1,2,\dots\}$ we set $n]=\{1,2,\dots,n\}$. Let $B_{n}=\{0,1\}^{n}$ be the set of binary $n$-tuples
+$\overline{b}=(\overline{b}_{1},\,\overline{b}_{2},\,\dots,\,\overline{b}_{n}),\ \overline{b}_{i}\in\{0,\,1\}\,.$
+We denote by
+$\|\overline{b}\|=\sum_{i=1}^{n}\overline{b}_{i}$}
+the number of $1$s in $\overline{b}$.
+For a finite set $A$ we denote by $|A|$ ($\in\mathbb{N}_{0}=\{0,1,\dots\}$) the number of its elements.
+For example, $|B_{n}|=2^{n}$.
+For a tuple $\overline{s}$ we denote by $|\overline{s}|$ its length, the number of terms in $\overline{s}$.
 By $\mathbb{Q}$ we denote the set of fractions.
-Let $n\in\mathbb{N}$ and $p=\frac{k}{l}\in(0,1]\cap\mathbb{Q}$, with $k,l\in\mathbb{N}$, $k\leq l$ and coprime $k,l$.
-We always assume that $p>0$.
-The $(p,n)$-proofreader is the pair $(P,\overline{s})$ in which
-$P=P(\overline{a},\,\overline{b})\colon B_{n}\times B_{n}\to B_{n},\ P(\overline{a},\,\overline{b})=(a_{1}b_{1},\,a_{2}b_{2},\,\dots,\,a_{n}b_{n})\,,$
-and
-$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N}),\ N=l^{n}\,\text{ and }\,\overline{s}_{j}\in B_{n}\,,$
-is a uniquely determined $N$-tuple defined as follows. We take the map $f\colon[l]\to\{0,1\}$, $f(x)=1$ for $x\in[k]$ and $f(x)=0$ for $x\in[l]\setminus[k]$, and extend it to
-$f\colon[l]^{n}\to B_{n},\ f((a_{1},\,a_{2},\,\dots,\,a_{n}))=(f(a_{1}),\,f(a_{2}),\,\dots,\,f(a_{n}))\,.$
-Then we take the lexicographic ordering
-$(\overline{u}_{1},\,\overline{u}_{2},\,\dots,\,\overline{u}_{N})$
-of $[l]^{n}$, so that $\overline{u}_{1}=(1,1,\dots,1)$, $\overline{u}_{2}=(1,1,\dots,1,2)$, $\dots$, $\overline{u}_{N}=(l,l,\dots,l)$.
-Any ordering of $[l]^{n}$ would do, but we want that $(p,n)$-proofreaders be unique objects.
+Let $p=k/l$ be in $(0,1]\cap\mathbb{Q}$, with $\in\mathbb{N}, klk\leq l$ and coprime $k,l$, and let $n\in\mathbb{N}$. We define the key discrete structure of our article.
+Definition 1.1 ($(p,n)$-proofreaders)
+$(p,n)$-proofreader $\mathcal{P}$ is the pair $\mathcal{P}=(P,\overline{s})$ with the following components which are uniquely determined by $p$ and $n$.
+  1. 1.
+    $P=P(\overline{a},\overline{b})\colon B_{n}\times B_{n}\to B_{n}$ is a map given by
+    $P(\overline{a},\,\overline{b})=(\overline{a}_{1}\overline{b}_{1},\,\overline{a}_{2}\overline{b}_{2},\,\dots,\,\overline{a}_{n}\overline{b}_{n})\,.$
+  2. 2.
+    The $N$-tuple of $n$-tuples
+    $\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\,$,
+    where $\overline{s}_{j}\in B_{n}$ and $N=l^{n}$, is defined as follows.
+We begin with the map $f\colon[l]\to\{0,1\}$ given by $f(x)=1$ for $x\in[k]$ and $f(x)=0$ for $x\in[l]\setminus[k]$.
+We extend it to the map $f\colon[l]^{n}\to B_{n}$ by
+    $f(a_{1},\,a_{2},\,\dots,\,a_{n})=(f(a_{1}),\,f(a_{2}),\,\dots,\,f(a_{n}))\,.$
+    We take the lexicographic ordering $\overline{u}_{1}$, $\overline{u}_{2}$, $\dots$, $\overline{u}_{N}$ of $l]^{n}$, so that $\overline{u}_{1}=(1,1,\dots,1)$, $\overline{u}_{2}=(1,1,\dots,1,2)$, $\dots$, $\overline{u}_{N}=(l,l,\dots,l)$.
 Finally, we set
     $\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})=(f(\overline{u}_{1}),\,f(\overline{u}_{2}),\,\dots,\,f(\overline{u}_{N}))\,.$
-Let $(P,\overline{s})$ be the $(p,n)$-proofreader and $\overline{t}\in B_{n}$ be a text, in which the errors are positions of $1$s.
-Then for $j\in[N]=[l^{n}]$, where $l$ is the denominator of $p$, the proofreader finds in the text $\overline{t}$ in the $j$-th run $\overline{s}_{j}$ exactly $\|P(\overline{s}_{j},\overline{t})\|$ errors.
-We show that proofreaders work in the desired way.
+    We call the $N$-tuple $\overline{s}$ the runs of $\mathcal{P}$.
+Any ordering of $l]^{n}$ would do. We chose the lexicographic one just concreteness sake.
+We show that the proofreaders introduced in Definition 1.1 work according to the parameters $p$ and $n$.
+The parameter $p$ is the efficiency (probability of error detection) of the proofreader, and the parameter $n$ is the length of the text submitted for proofreading.
+$\mathcal{P}=(P,\overline{s})$ works in $N$ runs $\overline{s}_{j}$, $j\in[N]$.
+If $(\overline{s}_{j})_{i}=1$ then in the $j$-th run $\mathcal{P}$ pays attention to the $i$-th position in the text, and if $(\overline{s}_{j})_{i}=0$ then $\mathcal{P}$ ignores it.
+So if a text $\overline{t}\in B_{n}$ (in which $1$s are errors) is submitted for proofreading, then $\mathcal{P}$ in the run $\overline{s}_{j}$ detects
+$\|P(\overline{s}_{j},\,\overline{t})\|\ \ (\in\mathbb{N}_{0})$
+errors in $\overline{t}$.
+The next proposition shows that for every set $I\subset[n]$ of positions in the text, $\mathcal{P}$ pays attention to all positions in II in exactly $p^{|I|}\cdot N$ of the $N$ runs.
 Proposition 1.2
-Let $(P,\overline{s})$, where
-$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\ \ (\in B_{n}^{N})\,,$
-be the $(p,n)$-proofreader. Then for every set $I\subset[n]$ we have
-$N^{-1}\cdot\big{|}\big{\{}j\in[N]\colon\;i\in I\Rightarrow s_{j,\,i}=1\big{\}}\big{|}=p^{|I|}\,.$
+Let $\overline{s}=(\overline{s}_{1},\overline{s}_{2},\dots,\overline{s}_{N})$ be the runs of the $(p,n)$-proofreader. Then for every set $I\subset[n]$ we have
+$N^{-1}\cdot\big|\big\{j\in[N]\colon\;\text{$(\overline{s}_{j})_{i}=1$ for every $i\in I$}\}\big|=p^{|I|}\,.$
 Proof.
-Let $\overline{s}$ and $I$ be as stated.
-It is easy to see from the definition of $\overline{s}$ that the number of $j\in[N]$ such that $s_{j,i}=1$ for every $i\in I$ is $k^{|I|}\cdot l^{n-|I|}$.
-But then
-$N^{-1}\cdot k^{|I|}\cdot l^{n-|I|}=\frac{k^{|I|}l^{n-|I|}}{l^{n}}=\bigg{(}\frac{k}{l}\bigg{)}^{|I|}=p^{|I|}\,.$
+Let $\overline{s}$, $p$, $n$, and $I$ be as stated.
+It is easy to see from the definition of $\overline{s}$ that the number of indices $j\in[N]$ such that $(\overline{s}_{j})_{i}=1$ for every $i\in I$ equals
+$k^{|I|}\cdot l^{n-|I|}$
+because it is the number of words of length $n$ over the alphabet $[l]$ that have at every position $i\in I$ a letter from $[k]$ ($\subset[l]$) . But then
+$N^{-1}\cdot k^{|I|}\cdot l^{n-|I|}=\frac{k^{|I|}\cdot l^{n-|I|}}{l^{n}}=\bigg(\frac{k}{l}\bigg)^{|I|}=p^{|I|}\,.$
 $\Box$
-Proofreading of a text is usually done only once, or a few times. We think of the $N$ runs $\overline{s}_{j}$ as of performances of the proofreader in possible worlds ([9]).
-We need two independent proofreaders.
-Let $n\in\mathbb{N}$, $p,q\in(0,1]\cap\mathbb{Q}$ and let $(P,\overline{s})$ and $(P,\overline{u})$ be the $(p,n)$-proofreader and the $(q,n)$-proofreader, respectively, with $|\overline{s}|=M$ and $|\overline{u}|=N$.
-Two (independent) $(p,q,n)$-proofreaders is the pair $(P,\overline{S})$ such that
+For Pólya’s argument we need two independent proofreaders. Suppose that $p,q\in(0,1]\cap\mathbb{Q}$ and that $n\in\mathbb{N}$.
+Definition 1.3 (two $(p,q,n)$-proofreaders)
+Two $(p,q,n)$-proofreaders is the pair $\mathcal{R}=(P,\overline{S})$ with the following components which are uniquely determined by $p$, $q$ and $n$.
+  1. 1.
+    The map $P$ is the same as in Definition 1.1.
+  2. 2.
+    We construct the $MN$-tuple $\overline{S}$ of pairs of $n$-tuples as follows.
+Let $(P,\overline{s})$ be the $(p,n)$-proofreader, $(P,\overline{u})$ be the $(q,n)$-proofreader, and let $M=|\overline{s}|$ and $N=|\overline{u}|$.
+We set
     $\overline{S}$ $=$ $(\overline{S}_{1},\,\overline{S}_{2},\,\dots,\,\overline{S}_{MN})$
-$=$ $((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{1},\,\overline{u}_{N}),\,(\overline{s}_{2},\,\overline{u}_{1}),\,(\overline{s}_{2},\,\overline{u}_{2}),\,\dots$
-$\dots,\,(\overline{s}_{2},\,\overline{u}_{N}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{1}),\,(\overline{s}_{M},\,\overline{u}_{2})\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))$
-($\in(B_{n}^{2})^{MN}$). Again, any ordering of the $MN$ pairs $(\overline{s}_{j},\overline{u}_{j^{\prime}})$ would do, but we want that pairs of proofreaders be unique. For any $j\in[MN]$ we write
+    $=$ $((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{1},\,\overline{u}_{N}),\,(\overline{s}_{2},\,\overline{u}_{1}),\,(\overline{s}_{2},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{2},\,\overline{u}_{N})\,$,
+    $\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{1}),\,(\overline{s}_{M},\,\overline{u}_{2}),\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\ \ (\in(B_{n}\times B_{n})^{MN})\,.$
+    We call the $MN$-tuple $\overline{S}$ the runs of $\mathcal{R}$.
+Any ordering of the $MN$ pairs $(\overline{s}_{j},\overline{u}_{j^{\prime}})$ would do, but we again want to have a uniquely defined object. For any $j\in[MN]$ we write
 $\overline{S}_{j}=\big(\overline{s}_{v_{j}},\,\overline{u}_{w_{j}}\big)\,.$
-We show that two proofreaders work in the desired way.
+In the next proposition, we show that for every set $I\subset[n]$ of positions in the text submitted for proofreading, the two proofreaders pay attention to these positions mutually independently.
 Proposition 1.4
-Let $(P,\overline{S})$, where
-$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
-be two $(p,q,n)$-proofreaders. The following holds.
+Let $\overline{S}=((\overline{s}_{1},\overline{u}_{1}),(\overline{s}_{1},\overline{u}_{2}),\dots,\dots,(\overline{s}_{M},\,\overline{u}_{N}))$ be the runs of two $(p,q,n)$-proofreaders.
+The following holds.
   1. 1.
     For every set $I\subset[n]$ we have
-    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow s_{v_{j},\,i}=1\big{\}}\big{|}=p^{|I|}\,.}$
+    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\text{$(\overline{s}_{v_{j}})_{i}=1$ for every $i\in I$}\big\}\big|=p^{|I|}\,.$}
   2. 2.
     For every set $I\subset[n]$ we have
-    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow u_{w_{j},\,i}=1\big{\}}\big{|}=q^{|I|}\,.}$
+    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\text{$(\overline{u}_{w_{j}})_{i}=1$ for every $i\in I$}\big\}\big|=q^{|I|}\,.$}
   3. 3.
     For every set $I\subset[n]$ we have
-    ${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;i\in I\Rightarrow s_{v_{j},\,i}=u_{w_{j},\,i}=1\big{\}}\big{|}=(pq)^{|I|}\,.}$
-Proof.
-Let $\overline{S}$ and $I$ be as stated.
-1.
-Let $r$ be the number of $j\in[M]$ such that $s_{j,i}=1$ for every $i\in I$.
-By Proposition 1.1, $\frac{r}{M}=p^{|I|}$.
-Thus the stated product equals
-${\textstyle\frac{1}{MN}\cdot rN=\frac{r}{M}=p^{|I|}\,.}$
-2. The argument is symmetric, with the number $s$ of $j^{\prime}\in[N]$ such that $u_{j^{\prime},i}=1$ for every $i\in I$. 3. Now the stated product equals
-${\textstyle\frac{1}{MN}\cdot rs=\frac{r}{M}\cdot\frac{s}{N}=p^{|I|}\cdot q^{|I|}=(pq)^{|I|}\,.}$
+    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\text{$(\overline{s}_{v_{j}})_{i}=(\overline{u}_{w_{j}})_{i}=1$ for every $i\in I$}\big\}\big|=(pq)^{|I|}\,.$}
+Proof. Let $\overline{S}$ and $I$ be as stated.
+1. Let $r$ be the number of indices $j\in[M]$ such that $(\overline{s}_{j})_{i}=1$ for every $i\in I$. By Proposition 1.2, $\frac{r}{M}=p^{|I|}$. Thus the displayed product equals
+$\frac{1}{MN}\cdot rN=\frac{r}{M}=p^{|I|}\,.$}
+2. The argument is symmetric, with the number $s$ of indices $j\in[N]$ such that $(\overline{u}_{j})_{i}=1$ for every $i\in I$.
+3. Now the displayed product equals
+$\frac{1}{MN}\cdot rs=\frac{r}{M}\cdot\frac{s}{N}=p^{|I|}\cdot q^{|I|}=(pq)^{|I|}\,.$}
 $\Box$
-We state our main result, a combinatorial version of Pólya’s trick.
+The following theorem is one of our main results.
 Theorem 1.5
-Let $(P,\overline{S})$, where
+Let $\mathcal{R}=(P,\overline{S})$, where
 $\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
 be two $(p,q,n)$-proofreaders and let
-${\textstyle c(p,\,q)=\frac{1}{2}\big{(}\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big{)}^{-1}\,.}$
-Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have for every number $a\in[0,m]$ that
+$c(p,\,q)=\frac{1}{2}\big(\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big)^{-1}\ \ (\in(0,\,\frac{1}{8}])\,.$}
+Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors and for every real number $a\in[0,m]$ we have
 ${\textstyle\frac{1}{MN}}\cdot\big|\big\{j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq a\big\}\big|\geq$
-${\textstyle\geq 1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})\,.}$
-For $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$ we define the fraction as +$\infty$+ ∞.
-Pólya’s estimate has distance at most $a$ from the number of errors for at least the fraction $1-\dots$ of pairs $(\overline{s}_{v_{j}},\overline{u}_{w_{j}})$ of runs of the two proofreaders.
-Next, in one extreme regime of Theorem 1.3 we minimize the bound $a$ on the distance.
-In the other extreme we maximize the fraction of pairs of runs for which the distance is still reasonably small.
-Deduction of both corollaries is immediate.
+${\textstyle\geq 1-6\exp\big(-\frac{2c(p,\,q)^{2}a^{2}}{m}\big)\,,}$
+where for $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$ we define the fraction $\frac{\cdots}{0}$ as $\infty$.
+The theorem says that the distance between Pólya’s approximation and the actual number $m$ of errors is at most $a$ for at least the displayed fraction 11-$\dots$ of the $MN$ runs of $\mathcal{R}$.
+The theorem has two corollaries.
+We minimize the distance of the approximation from the actual number of errors.
+Then we maximize the fraction of runs for which the distance is still reasonably small.
+The proofs are immediate, and we omit them.
 Corollary 1.6
-In the setup of Theorem 1.3, if $\omega(m)\colon\mathbb{N}\to\mathbb{N}$ is any function such that $\omega(m)\leq\sqrt{m}$ and $\omega(m)\to+\infty$, then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
+In the setup of Theorem 1.5, if $\omega\colon\mathbb{N}\to\mathbb{N}$ is a function such that
+[…]
+then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
 ${\textstyle\frac{1}{MN}}\cdot\big|\big\{j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq\omega(m)\sqrt{m}\big\}\big|\geq$
-$\geq 1-6\exp(-2c(p,\,q)^{2}\omega(m)^{2})\to 1\ \ (m\to+\infty)\,.$
+$\geq 1-6\exp(-2c(p,\,q)^{2}\omega(m)^{2})\to 1\ \ (m\to\infty)\,.$
 Corollary 1.7
@@
 ${\textstyle\frac{1}{MN}}\cdot\big|\big\{j\in[MN]\colon\;\left|\frac{\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|}{\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|}-m\right|\leq\varepsilon m\big\}\big|\geq$
-$\geq 1-6\exp(-2\varepsilon^{2}c(p,\,q)^{2}m)\to 1\ \ (m\to+\infty)\,.$
-We prove Theorem 1.3 next.
+$\geq 1-6\exp(-2\varepsilon^{2}c(p,\,q)^{2}m)\to 1\ \ (m\to\infty)\,.$
+We prove Theorem 1.5 in the next section.
 2 Proof of Theorem 1.5
-We deduce Theorem 1.3 from two propositions.
+We deduce Theorem 1.5 from the next proposition.
 Proposition 2.1
@@
 $\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
-are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
-${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\big{|}\|P(\overline{s}_{v_{j}},\,\overline{t})\|-pm\big{|}\leq a\big{\}}\big{|}\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
-An analogous bound holds when $\overline{s}_{v_{j}}$ replaced by $\overline{u}_{w_{j}}$, and $pm$ by $qm$.
-Proposition 2.2
-Suppose that $(P,\overline{S})$, where
+are the two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors the following holds.
+  1. 1.
+    We have
+    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\big|\|P(\overline{s}_{v_{j}},\,\overline{t})\|-pm\big|\leq a\big\}\big|\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
+  2. 2.
+    We have
+    $\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\big|\|P(\overline{u}_{w_{j}},\,\overline{t})\|-qm\big|\leq a\big\}\big|\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
+  3. 3.
+    We have
+    ${\textstyle\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\big|\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm\big|\leq a\big\}\big|}$
+    ${\textstyle\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
+We prove it at the end of this section. Now we deduce Theorem 1.5.
+Deduction of Theorem 1.5 from Proposition 2.1. Suppose that $(P,\overline{S})$, where
 $\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
-are two $(p,q,n)$-proofreaders and that $a>0$ is a real number. Then for every text $\overline{t}\in B_{n}$ with $\|\overline{t}\|=m\geq 1$ errors we have
-${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\big{|}\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm\big{|}\leq a\big{\}}\big{|}\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
-We prove both propositions at the conclusion of this section. Now we deduce from them Theorem 1.3.
-Deduction of Theorem 1.3 from Propositions 2.1 and 2.2. Suppose that $(P,\overline{S})$, where
-$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,,$
-are two $(p,q,n)$-proofreaders, that $\overline{t}\in B_{n}$ is a text with $\|\overline{t}\|=m\geq 1$ errors and that $a\in[0,m]$. We set
-${\textstyle b=c(p,\,q)\cdot a=\frac{1}{2}\big{(}\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big{)}^{-1}\cdot a\ \ (\leq\frac{pqm}{2})\,.}$
-For $j\in[MN]$, let $A$ be the set of $j$ such that $|\,\|P(\overline{s}_{v_{j}},\overline{t})\|-pm|\leq b$, $B$ be the set of $j$ such that $|\,\|P(\overline{u}_{w_{j}},\overline{t})\|-qm|\leq b$, $C$ be the set of $j$ such that
+are two $(p,q,n)$-proofreaders, that $\overline{t}\in B_{n}$ is a text with $\|\overline{t}\|=m\geq 1$ errors and that $a\in[0,m]$ is a real number. We set
+$b=c(p,\,q)\cdot a=\frac{1}{2}\big(\frac{1}{p}+\frac{1}{q}+\frac{2}{pq}\big)^{-1}\cdot a\ \ (\leq\frac{pqm}{2})\,.$}
+Let $A$ be the set of $j$ ($\in[MN])$ such that $|\,\|P(\overline{s}_{v_{j}},\overline{t})\|-pm|\leq b$, $B$ be the set of $j$ such that $|\,\|P(\overline{u}_{w_{j}},\overline{t})\|-qm|\leq b$, $C$ be the set of $j$ such that
 $|\,\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|-pqm|\leq b$
 and, finally, let $D$ be the set we are interested in, the set of $j$ such that
-$|R-m|:=\big{|}\,\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|\cdot\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|^{-1}-m\big{|}\leq a$
-($\|\dots\|\cdot\|\dots\|\cdot\|\dots\|^{-1}=+\infty$ if $\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=0$). Thus if $j\in A$, resp. $j\in B$, then
-[…]
+$|R-m|:=\big|\,\|P(\overline{s}_{v_{j}},\,\overline{t})\|\cdot\|P(\overline{u}_{w_{j}},\,\overline{t})\|\cdot\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|^{-1}-m\big|\leq a\,$,
+where $\|\dots\|\cdot\|\dots\|\cdot\|\dots\|^{-1}=+\infty$ if $\|P(\overline{s}_{v_{j}},P(\overline{u}_{w_{j}},\overline{t}))\|=0$. Thus if $j\in A$, resp. $j\in B$, then
+$\text{$\|P(\overline{s}_{v_{j}},\,\overline{t})\|=(1+\delta_{1})pm$, resp. $\|P(\overline{u}_{w_{j}},\,\overline{t})\|=(1+\delta_{2})qm$}\,$,
 where $|\delta_{1}|\leq\frac{b}{pm}$, resp. $|\delta_{2}|\leq\frac{b}{qm}$. Similarly, if $j\in C$ then
-${\textstyle\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=(1+\delta_{3})pqm,\ |\delta_{3}|\leq\frac{b}{pqm}\leq\frac{1}{2}\,.}$
+$\|P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t}))\|=(1+\delta_{3})pqm,\ |\delta_{3}|\leq\frac{b}{pqm}\leq\frac{1}{2}\,.$}
 If $j\in A\cap B\cap C$ then
@@
 $\leq$ $m\left(2|\delta_{3}|+2(|\delta_{1}|+|\delta_{2}|+|\delta_{1}|\cdot|\delta_{2}|)\right)$
-$\leq$ $2b{\textstyle\big{(}\frac{1}{pq}+\frac{1}{p}+\frac{1}{q}+\frac{1}{pq}\big{)}=a\,,}$
-and hence $D\supset A\cap B\cap C$. By Propositions 2.1 and 2.2,
-${\textstyle\frac{1}{MN}\cdot|D|\geq 1-\frac{|A^{c}|+|B^{c}|+|C^{c}|}{MN}\geq 1-6\exp(-\frac{2b^{2}}{m})=1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})\,.}$
-$\Box$
-We develop tools for proofs of Propositions 2.1 and 2.2. Let $(P,\overline{S})$, where
+$\leq$ $2b{\textstyle\big(\frac{1}{pq}+\frac{1}{p}+\frac{1}{q}+\frac{1}{pq}\big)=a}$
+and therefore $D\supset A\cap B\cap C$. By items 1–3 of Proposition 2.1,
+$\frac{1}{MN}\cdot|D|\geq 1-\frac{|A^{c}|+|B^{c}|+|C^{c}|}{MN}\geq 1-6\exp(-\frac{2b^{2}}{m})=1-6\exp(-\frac{2c(p,\,q)^{2}a^{2}}{m})$}
+(for $X\subset[MN]$ we write $X^{c}$ for the complement $MN]\setminus X$). $\Box$
+We develop tools for the proof of Proposition 2.1. Let $(P,\overline{S})$, where
 $\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
 be two $(p,q,n)$-proofreaders. We consider the finite uniform probability space
-$P=P(MN)=([MN],\,\mathcal{P}([MN]),\,\mathrm{Pr})\,,$
-in which $\mathrm{Pr}(\{j\})=\frac{1}{MN}$, and for $i\in[n]$ the three $n$-tuples of $P$-events
-$A_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=1\},\ A_{i}^{\prime}=\{j\in[MN]\colon\;u_{w_{j},i}=1\}$
+$\mathrm{PS}(MN)=([MN],\,\mathcal{P}([MN]),\,\mathrm{Pr})\,$,
+where $\mathrm{Pr}(\{j\})=\frac{1}{MN}$, and for $i\in[n]$ the three $n$-tuples of $\mathrm{PS}(MN)$-events
+$A_{i}=\{j\in[MN]\colon\;(\overline{s}_{v_{j}})_{i}=1\},\ A_{i}^{\prime}=\{j\in[MN]\colon\;(\overline{u}_{w_{j}})_{i}=1\}$
 and
-$C_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=u_{w_{j},i}=1\}\,.$
-By Propositions 1.1 and 1.2, each of these three $n$-tuples is independent.
-Further, $\mathrm{Pr}(A_{i})=p$, $\mathrm{Pr}(A_{i}^{\prime})=q$ and $\mathrm{Pr}(C_{i})=pq$ for every $i\in[n]$.
-Recall that $B^{c}$ denotes the complementary event to the event $B$ in a probability space.
+$C_{i}=\{j\in[MN]\colon\;(\overline{s}_{v_{j}})_{i}=(\overline{u}_{w_{j}})_{i}=1\}\,.$
+By items 1–3 of Proposition 1.4, each of these three $n$-tuples consists of independent events.
+Also, $\mathrm{Pr}(A_{i})=p$, $\mathrm{Pr}(A_{i}^{\prime})=q$ and $\mathrm{Pr}(C_{i})=pq$ for every $i\in[n]$.
+Recall that $B^{c}$ denotes the complementary event to the event $B$.
 For example,
-$A_{i}^{c}=[MN]\setminus A_{i}=\{j\in[MN]\colon\;s_{v_{j},i}=0\}\,.$
-Let $X\subset[n]$ be a nonempty set. For any set $Y\subset X$ we consider the three $P$-events
-${\textstyle A_{Y,X}=\bigcap_{i\in Y}A_{i}\cap\bigcap_{i\in X\setminus Y}A_{i}^{c},\ A_{Y,X}^{\prime}=\bigcap_{i\in Y}A_{i}^{\prime}\cap\bigcap_{i\in X\setminus Y}(A_{i}^{\prime})^{c}}$
+$A_{i}^{c}=[MN]\setminus A_{i}=\{j\in[MN]\colon\;(\overline{s}_{v_{j}})_{i}=0\}\,.$
+Let $X\subset[n]$ be a nonempty set. For any set $Y\subset X$ we consider the three $\mathrm{PS}(MN)$-events
+$A_{Y,X}=\bigcap_{i\in Y}A_{i}\cap\bigcap_{i\in X\setminus Y}A_{i}^{c},\ A_{Y,X}^{\prime}=\bigcap_{i\in Y}A_{i}^{\prime}\cap\bigcap_{i\in X\setminus Y}(A_{i}^{\prime})^{c}$}
 and
-${\textstyle C_{Y,X}=\bigcap_{i\in Y}C_{i}\cap\bigcap_{i\in X\setminus Y}C_{i}^{c}\,.}$
-We use the following key theorem which restates [2, Corollary A.1.7 on p. 325]. It is proven on [2, pp. 321–325].
+$C_{Y,X}=\bigcap_{i\in Y}C_{i}\cap\bigcap_{i\in X\setminus Y}C_{i}^{c}\,.$}
+We employ the following theorem which follows from [2, Corollary A.1.7 on p. 325] which is proven in [2, pp. 321–325].
 Theorem 2.2
-Let $n\in\mathbb{N}$, $X\subset[n]$ with $|X|=m\geq 1$ and let $a>0$ be a real number. With the above notation we have the bound
-${\textstyle\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.}$
-Two analogous bounds hold when $A_{Y,X}$ is replaced by $A_{Y,X}^{\prime}$, resp. $C_{Y,X}$, and $pm$ by $qm$, resp. $pqm$.
-We explain in detail this restatement.
+Let $n\in\mathbb{N}$, $X\subset[n]$ with $|X|=m\geq 1$ and let $a>0$ be a real number. With the above notation we have the following bounds.
+  1. 1.
+    $\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
+  2. 2.
+    $\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-qm\,|\leq a\end{subarray}}\mathrm{Pr}(A_{Y,X}^{\prime})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
+  3. 3.
+    $\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |\,|Y|-pqm\,|\leq a\end{subarray}}\mathrm{Pr}(C_{Y,X})\geq 1-2\exp(-\frac{2a^{2}}{m})\,.$}
+We explain in detail the deduction of this theorem from [2, Corollary A.1.7].
 Let $m\in\mathbb{N}$, $p\in[0,1]$, $X_{i}$ for $i\in[m]$ be $m$ independent two-valued real random variables in a probability space, with the common distribution $\mathrm{Pr}(X_{i}=1-p)=p$ and $\mathrm{Pr}(X_{i}=-p)=1-p$, and let
 $X_{0}=X_{1}+X_{2}+\dots+X_{m}\,.$
-The random variable $X_{0}$ has the centered binomial distribution $B(m,p)-pm$. The estimate [2, Corollary A.1.7], more precisely a particular case of it with all $p_{i}=p$, says the following.
+The random variable $X_{0}$ has the centered binomial distribution B(m,p)pmB(m,p)-pm. The estimate [2, Corollary A.1.7], more precisely a particular case of it with all $p_{i}=p$, says the following.
 Theorem 2.3 (Corollary A.1.7 in [2])
-For every real number $a>0$,
-${\textstyle\mathrm{Pr}(|X_{0}|>a)<\exp(-\frac{2a^{2}}{m})\,.}$
-What does the independence of random variables $X_{i}$, $i\in[m]$, which is assumed in Theorem 2.4, precisely mean?
+With the above notation we have for every real number $a>0$ that
+$\mathrm{Pr}(|X_{0}|>a)<\exp(-\frac{2a^{2}}{m})\,.$}
+What does the independence of random variables $X_{i}$, $i\in[m]$, which is needed in Theorem 2.3, precisely mean?
 Some events $C_{1}$, $C_{2}$, $\dots$, $C_{m}$ in a probability space are independent, if for every set $I\subset[m]$ we have
-${\textstyle\mathrm{Pr}\big{(}\bigcap_{i\in I}C_{i}\big{)}=\prod_{i\in I}\mathrm{Pr}(C_{i})\,.}$
-These events are strongly independent, if every $m$-tuple of events $D_{1}$, $D_{2}$, $\dots$, $D_{m}$, where $D_{i}\in\{C_{i},C_{i}^{c}\}$, is independent.
+$\mathrm{Pr}\big(\bigcap_{i\in I}C_{i}\big)=\prod_{i\in I}\mathrm{Pr}(C_{i})\,.$}
+These events are strongly independent, if every $m$-tuple of events $D_{1}$, $D_{2}$, $\dots$, $D_{m}$, where each $D_{i}\in\{C_{i},C_{i}^{c}\}$, is independent.
 A well known lemma, whose proof we leave for the interested reader as an exercise, says that any independent events are strongly independent.
 Discrete (i.e., with at most countable images) real random variables $Y_{1}$, $Y_{2}$, $\dots$, $Y_{m}$ are independent if for every $m$-tuple
-$\overline{a}=(a_{1},\,a_{2},\,\dots,\,a_{m})\in\mathbb{R}^{m}$
-the events $\mathrm{Pr}(Y_{i}=a_{i})$, $i\in[m]$, are independent.
-If the $Y_{i}$ are two-valued, the image of $Y_{i}$ being $\{a_{i},b_{i}\}$ with $a_{i}\neq b_{i}$, then, due to the mentioned lemma, the random variables $Y_{i}$ are independent if for any single $m$-tuple $\overline{c}\in\mathbb{R}^{m}$ with $c_{i}\in\{a_{i},b_{i}\}$ the events $\mathrm{Pr}(Y_{i}=c_{i})$, $i\in[m]$, are independent.
-We deduce Theorem 2.3 from Theorem 2.4.
-We consider only events $A_{Y,X}$, for events $A_{Y,X}^{\prime}$ and $C_{Y,X}$ we would argue similarly, relying on part 2 and 3 of Proposition 2.2.
-So let $(P,\overline{S})$, with
+$\overline{a}=(\overline{a}_{1},\,\overline{a}_{2},\,\dots,\,\overline{a}_{m})\in\mathbb{R}^{m}\,$,
+the events $\mathrm{Pr}(Y_{i}=\overline{a}_{i})$, $i\in[m]$, are independent.
+If the $Y_{i}$ are two-valued, the image of $Y_{i}$ being $\{a_{i},b_{i}\}$ with $a_{i}\neq b_{i}$, then, due to the mentioned lemma, the random variables $Y_{i}$ are independent if for any single $m$-tuple $\overline{c}\in\mathbb{R}^{m}$ with $\overline{c}_{i}\in\{a_{i},b_{i}\}$ the events $\mathrm{Pr}(Y_{i}=c_{i})$, $i\in[m]$, are independent.
+We deduce Theorem 2.2 from Theorem 2.3. We consider only item 1, the arguments for items 2 and 3 are similar. So let $(P,\overline{S})$, with
 $\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
@@
 We may assume that $X=[m]$.
-For $i\in[m]$ we define two-valued $P(MN)$-random variables $X_{i}$ by setting $X_{i}(j)=1-p$ if $j\in A_{i}$, and $X_{i}(j)=-p$ if $j\in[MN]\setminus A_{i}$.
-We set $X_{0}=\sum_{i=1}^{m}X_{i}$.
-Then for any subset $Y\subset X$ we have
+For $i\in[m]$ we define two-valued $\mathrm{PS}(MN)$-random variables $X_{i}$ by setting ($j\in[MN]$)
+$\text{$X_{i}(j)=1-p$ if $j\in A_{i}$, and $X_{i}(j)=-p$ if $j\in[MN]\setminus A_{i}$}\,.$
+We set $X_{0}=\sum_{i=1}^{m}X_{i}$. Then for any subset $Y\subset X$ we have
 $\mathrm{Pr}(A_{Y,X})=\mathrm{Pr}(X_{0}=|Y|-pm)\,.$
-By part 1 of Proposition 2.2 and the above discussion of independence, the random variables $X_{i}$, $i\in[m]$, are independent.
+By item 1 of Proposition 1.4 and the above discussion of independence, the random variables $X_{i}$, $i\in[m]$, are independent.
 Let $a>0$ be a real number.
-By applying Theorem 2.4 to the random variable $X_{0}$ we get the estimate in Theorem 2.3.
-We proceed to proofs of Propositions 2.1 and 2.2.
-Proof of Proposition 2.1. Suppose that $(P,\overline{S})$, where
+Applying Theorem 2.3 to the random variable $X_{0}$ we get the estimate in item 1 of Theorem 2.2.
+We proceed to the proof of Proposition 2.1.
+Proof of Proposition 2.1. We consider only item 1, the arguments for items 2 and 3 are similar. Suppose that $(P,\overline{S})$, where
 $\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
@@
 We may assume that the $1$s of $\overline{t}$ are on the first $m$ positions.
-We consider the $P(MN)$-events $A_{i}$, $i\in[m]$.
+We consider the $\mathrm{PS}(MN)$-events $A_{i}$, $i\in[m]$.
 Let $Y\subset[m]=X$.
 Then
-${\textstyle\frac{1}{MN}|\{j\in[MN]\colon\;\text{for $i\in[m]$ we have $s_{v_{j},i}=1$ iff $i\in Y$}\}|=\mathrm{Pr}(A_{Y,X})\,.}$
+$\frac{1}{MN}|\{j\in[MN]\colon\;\text{for $i\in[m]$ we have $(\overline{s}_{v_{j}})_{i}=1$ iff $i\in Y$}\}|=\mathrm{Pr}(A_{Y,X})\,.$}
 Thus for any $k\in\mathbb{N}_{0}$ we have
-${\textstyle\frac{1}{MN}\cdot\big{|}\big{\{}j\in[MN]\colon\;\|P(\overline{s}_{v_{j}},\,\overline{t})\|=k\big{\}}\big{|}=\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |Y|=k\end{subarray}}\mathrm{Pr}(A_{Y,X})\,.}$
-The stated estimate therefore follows from the first part of Theorem 2.3. The second part of Proposition 2.1 follows by a very similar argument from the second part of Theorem 2.3. $\Box$
-Proof of Proposition 2.2.
-Let $(P,\overline{S})$, $a>0$ and $\overline{t}$ be as in the previous proof.
-We consider the $P(MN)$-events $C_{i}$, $i\in[m]$, and argue as in the previous proof, relying on the third part of Theorem 2.3.
-$\Box$
+$\frac{1}{MN}\cdot\big|\big\{j\in[MN]\colon\;\|P(\overline{s}_{v_{j}},\,\overline{t})\|=k\big\}\big|=\sum_{\begin{subarray}{c}Y,\,Y\subset X\\ |Y|=k\end{subarray}}\mathrm{Pr}(A_{Y,X})\,.$}
+The stated estimate therefore follows from item 1 of Theorem 2.2. $\Box$
 3 Concluding remarks
-By [5, p.
-109], “The subject of probability is wholly free from real paradoxes.” This is not precise — not only books like [10] are devoted to paradoxes in probability theory, but some of these paradoxes are truly real.
-At their roots is the unrealistic nature, or even total absence, of physical experiments that should be modeled by the probabilistic model, or PM, at hand.
-In these paradoxical cases we often have a PM that does not model anything realistic.
-For example, take well known Bertrand’s paradox which asks about the probability that random chord in the unit circle in the plane has length at least $\sqrt{3}$ (the side of the inscribed equilateral triangle).
-The noun phrase “random chord” is problematic because it does not refer to anything clearly defined.
-In contrast, noun phrases like “random (finite!) graph” or “random subset (of a finite set)” are unproblematic; these concepts are easily modeled by finite uniform probability spaces and events in them, like the above $P(MN)$ and the $P(MN)$-events $A_{i}$, $A_{i}^{\prime}$ and $C_{i}$, $i\in[n]$.
-Mathematical chords have zero width and form an uncountable set.
-People try to resolve Bertrand’s paradox by rushing in defining PM on the set of mathematical chords.
-Instead, they should step back and consider physical chords with width $w>0$.
-A realistic and convincing physical discrete model, or DM, for selection of random physical chords should be devised.
-Only when such DM is understood, we can send $w$ to 00 and see what happens.
-This approach is implemented in [6].
-We hope to rework [6] in near future in a better form.
-In this article we approached Pólya’s proofreaders from the DM perspective. Returning to Proposition 1.1, we see that if $(P,\overline{s})$, where
-$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})\,,$
-is the $(p,n)$-proofreader, then in the uniform probability space
-$P(N)=([N],\,\mathcal{P}([N]),\,\mathrm{Pr})\,,$
-with $\mathrm{Pr}(\{j\})=\frac{1}{N}$, the $n$ $P(N)$-events
-$A_{i}=\{j\in[N]\colon\;s_{j,i}=1\},\ i\in[n]\,,$
-are independent and model random subset of $[n]$ with probability of an element equal to $p$.
-However, our construction requires $p$ ($\in(0,1]$) be a rational number.
-This is just as right because we cannot have a realistic physical experiment with probabilities like $\frac{1}{\sqrt{2}}$.
-Thus we conclude our article with the comment, which may come as a surprise to some, that the random graph $G(n,p)$ with probability of an edge equal to an irrational number $p$ is not really a realistic model.
+An interesting direction for further research is the following. Suppose that $(P,\overline{S})$, where
+$\overline{S}=((\overline{s}_{1},\,\overline{u}_{1}),\,(\overline{s}_{1},\,\overline{u}_{2}),\,\dots,\,\dots,\,(\overline{s}_{M},\,\overline{u}_{N}))\,$,
+are two $(p,q,n)$-proofreaders and that $\overline{t}\in B_{n}$ is a text. We do not have access to $\overline{t}$ and to $m=\|\overline{t}\|$, but the multiset of triples of values
+$V=\{(P(\overline{s}_{v_{j}},\,\overline{t}),\,P(\overline{u}_{w_{j}},\,\overline{t}),\,P(\overline{s}_{v_{j}},\,P(\overline{u}_{w_{j}},\,\overline{t})))\in\mathbb{N}_{0}^{3}\colon\;j\in[MN]\}$
+is known.
+Can we recover the number of errors $m$, or at least some bounds on $m$, from $V$?
+The parameter $n$ is of course known but we get two versions of this problem depending on whether $p$ and $q$ are known or not.
+Definitions 1.1 and 1.3 with accompanying Propositions 1.2 and 1.4 are probably more important results than Theorem 1.5 whose deduction from the well known Theorem 2.3 is relatively straightforward.
+The two definitions provide realistic models (RM) of Pólya’s proofreaders.
+Indeed, if
+$\overline{s}=(\overline{s}_{1},\,\overline{s}_{2},\,\dots,\,\overline{s}_{N})$
+are runs of a $(p,n)$-proofreader $(P,\overline{s})$ and $\overline{t}\in B_{n}$ is a text, then we can, at least in principle, physically go through all $N=l^{n}$ binary $n$-tuples $\overline{s}_{j}$ and compute all numbers $\|P(\overline{s}_{j},\overline{t})\|$ of errors the proofreader detects in the text during all $N$ runs.
+For this to be possible, it is crucial that $p\in\mathbb{Q}$.
+If $p=1/\sqrt{2}$, say, then no RM for a proofreader with efficiency $p$ exists.
+For irrational $p$ one has to be content with probabilistic models (PM), of random subsets containing (independently) a given element with probability $p$.
+Their construction is easy, but they cannot be physically realized in the same sense in which RM can.
+Some paradoxes in probability theory arise when a “practical” problem is modeled right away with a PM, while one should start carefully with a RM.
+The best example of this is the well known Bertrand paradox.
+This paradox begins as a “practical” problem: what is the probability that a random chord in the unit circle is longer than the length $\sqrt{3}$ of the side of the inscribed equilateral triangle.
+Solutions start with various PMs, which automatically assume dimensionless points and chords with zero width.
+This is a mistake, no such objects exist in reality.
+Instead, one should start with a RM where points have size $s>0$ and chords have width $w>0$.
+Once such RM is established, one can send $s,w\to 0$ and see what happens in the limit.
+This is the approach in [5], which we hope to rework and improve in the near future.